Vectors A and B lie in an xy-plane. A has magnitude 10.0 and an angle 109°; B has components Bx=-7.00 and By = -9.00. What isVectors A → and B →  lie in an xy-plane. A → has magnitude 10.0 and an angle 109°; B has components B x = -7.00 and B y = -9.00. What is 4A → .B → ? ( Your answer must be without decimals and maximum of 2% of error is accepted in your answer.)
Answer #1

and Ā= 10 (cos 109°? + sin 1097) Ã. B = [10 cos 109 * (-7)] + B=(-7 î–97) [losin 109 *(-9)] So, A. B = 22.79 + (-85.097) = –